OpenStudy (anonymous):

Evaluate the function f(p) = p2 + 3p + 1 for p = -2.

OpenStudy (anonymous):

By "p2," do you mean p^2 (p to the power of 2)?

OpenStudy (defiance):

Rewright it by filling p with -2.

OpenStudy (anonymous):

1 -1 -8 9 -9 are the answer choices...

OpenStudy (anonymous):

f(-2) = (-2)^2 +3(-2)+1 f(-2) = ???

OpenStudy (anonymous):

Now after seeing that, what do you think the answer is?

OpenStudy (anonymous):

it would be -9..

OpenStudy (anonymous):

try looking at this site http://tinyurl.com/2g9mqh

OpenStudy (anonymous):

By "p2" do you mean: \[p ^{2}\]

OpenStudy (defiance):

It's -9 either way. If you use exponent or "2" or times by 2... you still get the -4.

OpenStudy (anonymous):

thanks

OpenStudy (defiance):

Then you 3 x -2 = -6 -4 + -6 + 1 = -9